# An Accelerated Hybrid Low-Rank Augmented Lagrangian Method for Large-Scale Semidefinite Programming on GPU

### v1

@@ 4.3 Experiments for maximum stable set
Table 5 shows the significant performance advantage of GPU acceleration for larger-scale Maximum Stable Set problems on Hamming graphs. For small Hamming graphs, such as $H_{10,2}$ through $H_{13,2}$, HALLaR outperforms the cuHALLaR by factors of 0.088x to 0.596x, reflecting GPU parallelization overhead for modest problem sizes and the fact that HALLaR is not designed for small problems. A clear performance transition occurs at $H_{14,2}$, where cuHALLaR begins demonstrating superior performance with a 1.21x speedup. This advantage grows exponentially as problem size increases, reaching an impressive 50.1x speedup for $H_{23,2}$ with over 8 million vertices. All instances maintain a consistent optimal solution rank of 2, confirming the effectiveness of our low-rank approach. Notably, cuLoRADS encounters memory limitations (OOM) for problems larger than $H_{14,2}$, highlighting the superior memory efficiency of our implementation. For problems it could solve, cuLoRADS shows significantly longer runtimes and higher ranks than both our GPU and CPU implementations.
The GSET benchmark [75] highlights how GPU performance shows a strong correlation with problem structure: for dense graphs with high-rank solutions such as G58, G59, and G64, GPU acceleration yields 4-5x speedups. However, for sparse graphs with very low-rank solutions such as G11, G12, and G32, CPU implementation performs 30-100x faster. This suggests that GPU acceleration is most beneficial when the computational workload involves more complex matrix operations that can effectively utilize parallel processing. Interestingly, the solution ranks sometimes differ between implementations, with GPU occasionally finding different rank solutions than CPU. The performance correlation appears more strongly tied to edge density than vertex count, with the highest speedups observed on instances having both moderate-to-large size and high edge count. These findings indicate that problem-specific characteristics, rather than simply problem size, determine when GPU acceleration provides the most significant performance advantages.
Problem Instance HALLaR cuHALLaR Ratio
G1(800; 19,176) 98.25/98 45.19/98 2.17
G10(800; 19,176) 80.51/97 40.52/97 1.99
G11(800; 1,600) 0.14/2 4.23/2 0.03
G12(800; 1,600) 0.09/2 2.59/2 0.03
G14(800; 4,694) 45.37/73 214.73/74 0.21
G20(800; 4,672) 129.73/124 359.40/166 0.36
G43(1000; 9,990) 34.20/60 35.84/60 0.95
G51(1,000; 5,909) 167.03/141 273.84/128 0.61
G23(2,000; 19,990) 94.81/76 52.59/76 1.80
G31(2,000; 19,990) 122.07/77 58.23/76 2.10
G32(2,000; 4,000) 0.35/2 25.34/3 0.01
G34(2,000; 4,000) 2.00/2 25.29/3 0.08
G35(2,000; 11,778) 397.32/133 254.68/108 1.56
G41(2,000; 11,785) 343.84/143 286.82/139 1.20
G48(3,000; 6,000) 2.29/2 18.95/2 0.12
G55(5,000; 12,498) 152.03/45 81.11/45 1.87
G56(5,000; 12,498) 151.99/45 80.83/45 1.88
G57(5,000; 10,000) 8.71/2 42.55/2 0.20
G58(5,000; 29,570) 2,191.49/112 504.55/130 4.35
G59(5,000; 29,570) 2,167.27/111 539.28/121 4.02
G60(7,000; 17,148) 262.35/50 96.89/50 2.71
G62(7,000; 14,000) 7.80/3 25.18/3 0.31
G64(7,000; 41,459) 1,719.79/51 334.12/51 5.15
G66(9,000; 18,000) 11.31/3 27.91/3 0.41
G67(10,000; 20,000) 2.39/2 5.83/2 0.41
G72(10,000; 20,000) 2.41/2 5.81/2 0.41
G77(14,000; 28,000) 27.34/3 11.81/2 2.31
G81(20,000; 40,000) 53.85/3 64.20/3 0.84
Table 6: Runtimes (in seconds) for the Maximum Stable Set problem on GSET instances. Tolerances are set to $10^{-5}$.
@@ 4.4 Experiments for phase retrieval
Given $m$ pairs $\{(a_{i},b_{i})\}_{i=1}^{m}\subseteq\mathbb{C}^{n}\times\mathbb{R}_{+}$, consider the problem of finding a vector $x\in\mathbb{C}^{n}$ such that
@@
$\min_{X}\quad\left\{\mathrm{tr}(X)\quad:\quad\langle a_{i}a_{i}^{H},X\rangle=b_{i},\quad X\succeq 0,\quad X\in\mathbb{S}^{n}(\mathbb{C})\right\},$
@@
Problem Instance HALLaR cuHALLaR Ratio
@@
3,162,278; 37,947,336 Timed Out 583.782/14 N/A
Table 7: Runtimes (in seconds) for the Phase Retrieval problem. A relative tolerance of $\epsilon=10^{-5}$ is set and a time limit of 43200 seconds (12 hours) is given.
In Table 7 we observe significant performance improvements with cuHALLaR compared to the CPU HALLaR as problem sizes increase. For medium-sized instances with $n=31,623$, cuHALLaR achieves consistent speedups of 3.86-5.23x, completing calculations in approximately 29-31 seconds versus 119-153 seconds on CPU. For larger instances with $n=100,000$, the speedup increases dramatically to 29.24-46.78x, with cuHALLaR solving problems in just 8.6-10.4 seconds compared to 304-434 seconds on CPU. The largest reported instances of $n=316,228$ demonstrate impressive speedups of 15.05-22.58x, with GPU solution times around 29-34 seconds compared to 508-699 seconds on CPU. Most notably, cuHALLaR successfully solves the massive instance with $n=3,162,278$ in under 10 minutes, while the CPU implementation times out after 12 hours. It’s important to highlight that cuLoRADS was unable to handle these phase retrieval problems at all due to its reliance on the SDPA format, which becomes prohibitively expensive in both memory and computational requirements for problems of this scale. This further demonstrates the superior memory efficiency and scalability of our GPU implementation compared to both CPU-based methods and other GPU-accelerated SDP solvers.
@@ 5 Concluding Remarks
Through extensive numerical experiments across three problem domains, we demonstrate that cuHALLaR achieves remarkable performance improvements over both CPU-based implementations and competing GPU-accelerated solvers. For matrix completion problems, cuHALLaR consistently outperforms the state-of-the-art GPU solver cuLoRADS by factors of 2x to 25x, while finding solutions with significantly lower ranks. Most impressively, cuHALLaR successfully solves instances with more than 300 million constraints in under 3 minutes, while the CPU implementation, HALLaR, fails to converge within 12 hours.
For maximum stable set problems, our results reveal that cuHALLaR’s performance advantage scales dramatically with problem size and graph density. We achieve speedups of 50x for the largest Hamming graph problems and 135x for million-node AG-Monien instances. Similarly, for phase retrieval problems, cuHALLaR delivers speed improvements exceeding 74x, solving massive instances with over 3 million dimensions in under 10 minutes—problems that remain intractable for CPU-based approaches.
Future work. While cuHALLaR demonstrates exceptional performance across the problems tested, certain structural characteristics of SDPs may impact its efficiency. Problems with inherently high-rank solutions or those requiring extremely high precision may benefit less from our approach. Additionally, our current implementation focuses on single-GPU execution and double-precision arithmetic. Future research directions include developing optimized multi-GPU implementations in C/C++ to further enhance scalability, exploring mixed-precision computation techniques to balance accuracy and performance, and extending our methodology to a broader class of SDP problems. These developments would further advance the approaches introduced in [47] and establish GPU-accelerated low-rank methods as the standard for solving large-scale SDPs in practical applications.
@@ Acknowledgments
This research was supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at Georgia Tech, Atlanta, Georgia, USA.
@@ Appendix A ADAP-AIPP
This section presents the ADAP-AIPP method first developed in [47, 60]. HALLaR uses ADAP-AIPP in its step 4 to find an approximate stationary point (according to the criteria in (6)) of a nonconvex problem of the form
$\min_{U}\quad\left\{g(U)=\mathcal{L}_{\beta}(UU^{\top};p)\quad:\quad U\in B_{1}^{s}\right\}$ (18)
where $\mathcal{L}_{\beta}(X;p)$ is as in (5) and $B_{1}^{s}$ is as in (2).
Given an initial point $\underline{W}\in B_{1}^{s}$ and a tolerance $\rho>0$, the goal of ADAP-AIPP is to find a a triple $(\overline{W};\overline{R},\rho)$ that satisfies
$\overline{R}\in\nabla g(\overline{W})+N_{B^{s}_{1}}(\overline{W}),\quad\|\overline{R}\|_{F}\leq\rho,\quad g(\overline{W})\leq g(\underline{W}).$ (19)
@@
Input: a function $g$ as in (18), an initial point $\underline{W}\in B_{1}^{s}$, an initial prox stepsize $\lambda_{0}>0$, and a tolerance $\rho>0$.
@@
    $\psi_{s}$ $={\lambda}g+\frac{1}{2}\|\cdot-W_{j-1}\|_{F}^{2},\quad\psi_{n}={\lambda}\delta_{B^{s}_{1}};$ (22)
@@
    ADAP-AIPP is a double looped algorithm, whose outer iterations are indexed by $j$. During its $j$-th iteration, ADAP-AIPP attempts to solve a proximal subproblem of the form $\min_{U\in B_{1}^{s}}\lambda g(U)+0.5\|\cdot-W_{j-1}\|^{2}_{F}$, where $\lambda$ is a prox stepsize and $W_{j-1}$ is the current prox center. If a proximal subproblem cannot suitably be solved, ADAP-AIPP reduces $\lambda$ and tries to solve the new subproblem.
@@ A.1 ADAP-FISTA
Input: a function $\psi_{s}$ as in (25), an initial point $x_{0}\in B_{1}^{s}$ and scalars $\mu>0$ and $L_{0}>\mu$.
@@
    $y_{i+1}$ $\;\;=\;\;\arg\min_{u\in B_{1}^{s}}\Bigl{\{}\ell_{\psi_{s}}(u;\tilde{x}_{i})+\tfrac{L_{i+1}}{2}\|u-\tilde{x}_{i}\|^{2}\Bigr{\}};$ (27)
@@
    $x_{i+1}$ $=\frac{1}{\tau_{i+1}}\bigl{[}\mu\,a_{i}\,y_{i+1}+\tau_{i}\,x_{i}-a_{i}\,s_{i+1}\bigr{]};$ (31)
@@
    $\|v_{i+1}\|\leq\sigma\|y_{i+1}-x_{0}\|$ (34)
    holds then stop with success and output $(y,v,L):=(y_{i+1},v_{i+1},L_{i+1})$; otherwise, $i\leftarrow i+1$ and go to step 1.

### v2

@@ 5.3 Experiments for maximum stable set
Table 4 compares HALLaR, cuHALLaR, and cuLoRADS on small Hamming graph instances, with the HALLaR and cuHALLaR benchmarks conducted using on-the-fly problem generation. For small graphs, such as $H_{10,2}$ through $H_{13,2}$, the CPU-based HALLaR outperforms cuHALLaR, reflecting the overhead of GPU parallelization on modest problem sizes. A clear performance transition occurs at $H_{14,2}$, where cuHALLaR begins to demonstrate superior performance with a 1.21x speedup. This advantage grows exponentially as problem size increases, reaching an impressive 50.1x speedup over HALLaR for the $H_{23,2}$ instance, which has over 8 million vertices and 96 million edges. Notably, cuLoRADS encounters out-of-memory limitations for problems larger than $H_{14,2}$, highlighting the superior memory efficiency of our implementations. For the problems cuLoRADS could solve, it took significantly longer and found solutions with higher ranks than both HALLaR and cuHALLaR.
The results for the GSET instances in Table 6, obtained via on-the-fly problem generation, show that performance is strongly correlated with graph structure. For sparse graphs that admit low-rank solutions (e.g., G11, G12, and G32), the CPU-based HALLaR is 30–100x faster than cuHALLaR. Conversely, for dense graphs yielding high-rank solutions, such as G58, G59, and G64, cuHALLaR achieves a 4–5x speedup.
Problem Instance On-the-fly Generation HSLR SDPA
Graph($n$; $|E|$) HALLaR cuHALLaR Ratio cuHALLaR (HSLR) cuLoRADS (SDPA) Ratio
G1(800; 19,176) 98.25/98 45.19/98 2.17 53.63/97 340.66 6.35
G10(800; 19,176) 80.51/97 40.52/97 1.99 53.11/95 330.18 6.22
G11(800; 1,600) 0.14/2 4.23/2 0.03 7.29/2 7.78 1.07
G12(800; 1,600) 0.09/2 2.59/2 0.03 6.60/2 8.32 1.26
G14(800; 4,694) 45.37/73 214.73/74 0.21 211.04/55 484.41 2.30
G20(800; 4,672) 129.73/124 359.40/166 0.36 */3.0e-05 518.64 $0.2<$
G43(1000; 9,990) 34.20/60 35.84/60 0.95 43.27/58 383.83 8.87
G51(1,000; 5,909) 167.03/141 273.84/128 0.61 289.18/127 540.96 1.87
G23(2,000; 19,990) 94.81/76 52.59/76 1.80 74.39/76 766.91 10.31
G31(2,000; 19,990) 122.07/77 58.23/76 2.10 73.34/76 841.21 11.47
G32(2,000; 4,000) 0.35/2 25.34/3 0.01 40.89/2 13.45 0.33
G34(2,000; 4,000) 2.00/2 25.29/3 0.08 47.80/2 16.67 0.35
G35(2,000; 11,778) 397.32/133 254.68/108 1.56 689.21/142 14.55 0.02
G41(2,000; 11,785) 343.84/143 286.82/139 1.20 432.96/189 1527.71 3.53
G48(3,000; 6,000) 2.29/2 18.95/2 0.12 40.62/2 42.57 1.05
G55(5,000; 12,498) 152.03/45 81.11/45 1.87 102.45/39 309.48 3.02
G56(5,000; 12,498) 151.99/45 80.83/45 1.88 102.66/39 299.69 2.92
G57(5,000; 10,000) 8.71/2 42.55/2 0.20 50.91/2 26.09 0.51
G58(5,000; 29,570) 2,191.49/112 504.55/130 4.35 388.23/67 3178.00 8.19
G59(5,000; 29,570) 2,167.27/111 539.28/121 4.02 386.05/65 3159.04 8.18
G60(7,000; 17,148) 262.35/50 96.89/50 2.71 145.95/52 455.19 3.12
G62(7,000; 14,000) 7.80/3 25.18/3 0.31 21.61/2 34.88 1.61
G64(7,000; 41,459) 1,719.79/51 334.12/51 5.15 489.03/72 1655.44 3.38
G66(9,000; 18,000) 11.31/3 27.91/3 0.41 73.86/2 37.28 0.50
G67(10,000; 20,000) 2.39/2 5.83/2 0.41 57.05/2 69.30 1.21
G72(10,000; 20,000) 2.41/2 5.81/2 0.41 65.19/2 33.56 0.51
G77(14,000; 28,000) 27.34/3 11.81/2 2.31 57.71/2 52.19 0.90
G81(20,000; 40,000) 53.85/3 64.20/3 0.84 66.28/2 71.37 1.08
Table 6: Runtimes (in seconds) for the Maximum Stable Set problem on GSET instances. Tolerances are set to $10^{-5}$. Time limit was set to 1 hour (3600 seconds). An entry marked with /N*/N means the corresponding method finds an approximate solution with relative accuracy strictly larger than the desired accuracy in which case NN expresses the maximum of the three relative accuracies in (31). Ratio in the fourth column is computed as HALLaR/ cuHALLaR. Ratio in the last column is computed as cuLoRADS/ cuHALLaR.
The last block of columns of Table 6 demonstrates a clear performance advantage of cuHALLaR (with HSLR format) over cuLoRADS, which utilizes standard SDPA format. Our method cuHALLaR achieves significant speedups over cuLoRADS on medium-to-large sized graphs with substantial edge density. For instance, on graphs G23 and G31, cuHALLaR achieves speedups of 10.3x and 11.5x, respectively. This trend continues for larger graphs like G58 and G59, where the speedup is consistently over 8x. However, the performance dynamic reverses for SDP instances where the underlying graph is very sparse. On graphs such as G32 and G34, where the number of edges is only twice the number of vertices, cuLoRADS is approximately three times faster than cuHALLaR. Similarly, for G57, G66, and G72, cuLoRADS is roughly two times faster than cuHALLaR. This performance dichotomy suggests that the architectural benefits of the HSLR format and the associated kernels in cuHALLaR are most pronounced when the computational workload per iteration is high. For highly sparse problems, the lower overhead of processing the simpler SDPA format in cuLoRADS proves more efficient, indicating that the relative performance of the solvers is strongly dependent on the underlying graph structure.
@@ 5.4 Experiments for phase retrieval
Given $m$ pairs $\{(a_{i},b_{i})\}_{i=1}^{m}\subseteq\mathbb{C}^{n}\times\mathbb{R}_{+}$, this subsection considers the problem of finding a vector $x\in\mathbb{C}^{n}$ such that
@@
$\min_{X\in\mathbb{S}^{n}(\mathbb{C})}\quad\left\{\mathrm{Tr}(X)\quad:\quad\langle a_{i}a_{i}^{H},X\rangle=b_{i},\quad X\succeq 0\right\},$ (34)
@@
The SDP relaxation of the phase retrieval problem is hard because the linear and adjoint operators require the utilization of a fast Fourier transform (FFT) subroutine. On the GPU, this subroutine can be handled through CUDA’s cuFFT package [51]. Moreover, the SDP instances are extremely dense, and thus storing the data is expensive. For these reasons, the cuHALLaR binary and cuLoRADS were not tested for this problem class. We only tested HALLaR and cuHALLaR using on-the-fly problem generation.
Problem Instance ($n$; $m$) HALLaR (on-the-fly) cuHALLaR (on-the-fly) Ratio
@@
3,162,278; 37,947,336 */4.1e-03 583.782/14 >74.0>74.0
Table 7: Runtimes (in seconds) and final rank for the Phase Retrieval problem. A relative tolerance of $\epsilon=10^{-5}$ is used, with a 12-hour time limit is set for HALLaR. An entry marked with /N*/N means the corresponding method finds an approximate solution with relative accuracy strictly larger than the desired accuracy in which case NN expresses the maximum of the three relative accuracies in (31). The “Ratio” column reports the ratio of HALLaR runtime to cuHALLaR runtime.
Table 7 shows that cuHALLaR achieves substantial acceleration over HALLaR for SDP instances of the phase retrieval problem. Interestingly, the speedup that cuHALLaR achieves over HALLaR is generally more significant as the size of the SDP instance increases. For instances with dimension n=31,623n=31,\!623, cuHALLaR achieves speedups of 3.86x to 5.23x over HALLaR. For larger instances with dimension $n=100,000$, the speedup that cuHALLaR achieves over HALLaR increases dramatically to 29.24x–46.78x. For phase retrieval instances considered with dimension $n=316,228$, cuHALLaR achieves speedups from 15.05x to 22.58x over HALLaR.
Most significantly, cuHALLaR can successfully solve a massive phase retrieval instance with dimension $n=3,162,278$ in under 10 minutes. In stark contrast, HALLaR fails to solve this instance within the 12-hour time limit. Hence, for this huge problem instance, cuHALLaR is at least 74 times faster than HALLaR.
@@ 6 Concluding Remarks
Through extensive numerical experiments across three problem domains, we demonstrate that cuHALLaR achieves remarkable performance improvements over both CPU-based implementations and competing GPU-accelerated solvers. For matrix completion problems, cuHALLaR consistently outperforms the state-of-the-art GPU solver cuLoRADS by factors of 2x to 25x, while finding solutions with significantly lower ranks. Most impressively, cuHALLaR successfully solves instances with approximately 260 million constraints in under 1 minute, while the CPU implementation, HALLaR, fails to converge within 12 hours.
For maximum stable set problems, our results reveal that cuHALLaR’s performance advantage scales dramatically with problem size and graph density. We achieve speedups of 50x for the largest Hamming graph problems and 135x for million-node AG-Monien instances. Similarly, for phase retrieval problems, cuHALLaR achieves speedups of 15x to 47x on large instances, and successfully solves the largest test case—with a matrix variable of size over 3 million—in under 10 minutes. In contrast, the CPU-based HALLaR times out after 12 hours on this same instance, rendering it effectively intractable.
Future work. While cuHALLaR demonstrates exceptional performance across the problems tested, certain structural characteristics of SDPs may impact its efficiency. Problems with inherently high-rank solutions or those requiring extremely high precision may benefit less from our approach. Additionally, our current implementation focuses on single-GPU execution and double-precision arithmetic. Future research directions include developing optimized multi-GPU implementations in C/C++ to further enhance scalability, exploring mixed-precision computation techniques to balance accuracy and performance, and extending our methodology to a broader class of SDP problems such as those with multi-block matrix variable and inequality constraints. These developments would further advance the approaches introduced in [50] and establish GPU-accelerated low-rank methods as the standard for solving large-scale SDPs in practical applications.
@@ Acknowledgments
This research was supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at Georgia Tech, Atlanta, Georgia, USA, along with allocation MTH250047 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants.
Appendix
@@ Appendix A ADAP-AIPP
This section presents the ADAP-AIPP method first developed in [50, 62]. HALLaR uses ADAP-AIPP in its step 5 to find an approximate stationary point (according to the criterion in (7)) of a nonconvex problem of the form
$\min_{U}\quad\left\{g(U)=\mathcal{L}_{\beta}(UU^{\top};p)\quad:\quad U\in B_{\tau}^{s}\right\}$ (35)
where $\mathcal{L}_{\beta}(X;p)$ is as in (6) and $B_{\tau}^{s}$ is as in (3).
Given an initial point $\underline{W}\in B_{\tau}^{s}$ and a tolerance $\rho>0$, the goal of ADAP-AIPP is to find a triple $(\overline{W};\overline{R},\rho)$ that satisfies
$\overline{R}\in\nabla g(\overline{W})+N_{B_{\tau}^{s}}(\overline{W}),\quad\|\overline{R}\|_{F}\leq\rho,\quad g(\overline{W})\leq g(\underline{W}).$ (36)
@@
Input: a function $g$ as in (35), an initial point $\underline{W}\in B_{\tau}^{s}$, an initial prox stepsize $\lambda_{0}>0$, and a tolerance $\rho>0$.
@@
    $\psi_{s}$ $={\lambda}g+\frac{1}{2}\|\cdot-W_{j-1}\|_{F}^{2},\quad\psi_{n}={\lambda}\delta_{B_{\tau}^{s}};$ (39)
@@
    ADAP-AIPP is a double-looped algorithm, whose outer iterations are indexed by $j$. During its $j$-th iteration, ADAP-AIPP attempts to solve a proximal subproblem of the form $\min_{U\in B_{\tau}^{s}}\lambda g(U)+0.5\|U-W_{j-1}\|^{2}_{F}$, where $\lambda$ is a prox stepsize and $W_{j-1}$ is the current prox center. If a proximal subproblem cannot suitably be solved, ADAP-AIPP reduces $\lambda$ and tries to solve the new subproblem.
@@ A.1 ADAP-FISTA
Input: a function $\psi_{s}$ as in (42), an initial point $x_{0}\in B_{\tau}^{s}$ and scalars $\mu>0$ and $L_{0}>\mu$.
@@
    $y_{i+1}$ $=\arg\min_{u\in B_{\tau}^{s}}\left\{\ell_{\psi_{s}}(u;\tilde{x}_{i})+\tfrac{L_{i+1}}{2}\|u-\tilde{x}_{i}\|^{2}\right\};$ (44)
@@
    $x_{i+1}$ $=\frac{1}{\tau_{i+1}}\left[\mu\,a_{i}\,y_{i+1}+\tau_{i}\,x_{i}-a_{i}\,s_{i+1}\right];$ (48)
@@
    $\|s_{i+1}\|\;\leq\;\sigma\,\|y_{i+1}-x_{0}\|$ (50)
    holds, then stop with success and output $(y,L)=(y_{i+1},L_{i+1})$; otherwise set $i\leftarrow i+1$ and return to step 1.